40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 11 18:11:46 2015, Koji, Summary, LSC, 3f modulation cancellation 
    Reply  Thu Feb 12 23:47:45 2015, Koji, Update, LSC, 3f modulation cancellation 
       Reply  Sat Feb 14 00:48:13 2015, Koji, Update, LSC, 3f modulation cancellation 
          Reply  Sat Feb 14 19:54:04 2015, Koji, Summary, LSC, 3f modulation cancellation beat_setup1.JPGbeat_setup2.JPGelectrical_setup.pdf
             Reply  Sat Feb 14 20:37:51 2015, Koji, Summary, LSC, 3f modulation cancellation beat_pd_response.pdfbeat_nominal.pdf3f_reduction.pdfbeat_3f_reduced.pdf
                Reply  Sat Feb 14 22:14:02 2015, Koji, Summary, LSC, [HOW TO] 3f modulation cancellation freq_gen_box.JPGdelay_line.JPGcable_spec.pdf
                   Reply  Sun Feb 15 16:20:44 2015, Koji, Summary, LSC, [ELOG LIST] 3f modulation cancellation 
                      Reply  Sun Feb 15 20:55:48 2015, rana, Summary, LSC, [ELOG LIST] 3f modulation cancellation 
                         Reply  Mon Feb 16 00:08:44 2015, Koji, Summary, LSC, [ELOG LIST] 3f modulation cancellation 
                   Reply  Mon Feb 16 01:45:12 2015, Koji, Summary, LSC, modulation depth analysis modulation_nominal.pdfmodulation_3f_reduced.pdf
Message ID: 11019     Entry time: Thu Feb 12 23:47:45 2015     In reply to: 11005     Reply to this: 11028
Author: Koji 
Type: Update 
Category: LSC 
Subject: 3f modulation cancellation 

- I built another beat setup on the PSL table at the South East side of the table.
- The main beam is not touched, no RF signal is touched, but recognize that I was present at the PSL table.
- The beat note is found. The 3rd order sideband was not seen so far.
- A PLL will be built tomorrow. The amplifier box Manasa made will be inspected tomorrow.

- One of the two beams from the picked-off beam from the main beam line was introduced to the beat setup.
(The other beam is used of for the beam pointing monitors)
There is another laser at that corner and the output from this beam is introduced into the beat setup.
The combined beam is introduced to PDA10CF (~150MHz BW).

- The matching of the beam there is poor. But without much effort I found the beat note.
  The PSL laser had 31.33 deg Xtal temp. When the beat was found, the aux laser had the Xtal temp of 40.88.

- I could observe the sidebands easily, with a narrower BW of the RF analizer I could see the sidebands up to the 2nd order.
  The 3rd order was not seen at all.

- The beat note had the amplitude of about -30dBm. One possibility is to amplify the signal. I wanted to use a spare channel
of the ALS/FOLL amplifier box. But it gave me rather attenuation than any amplification.
I'll look at the box tomorrow.

- Also the matching of two beams are not great. The PD also has clipping I guess. These will also be improved tomorrow

- Then the beat note will be locked at a certain frequency using PLL so that we can reduce the measurement BW more.

 

ELOG V3.1.3-